I need 9 tsp of baking powder but I have baking soda and 1/4 of that is equal to 1 tsp of powder.

How much baking soda is needed?
Answer #1

From the given data.

Required Baking Powder = 9 tsp

But as per the question, we have Baking Soda instead of Baking Powder.

1/4 tsp of Baking Soda = 1 tsp of Baking Powder

We require 9 tsp of Baking Powder, Hence Multiply 9 on both sides

9*(1/4) tsp baking soda = 9 * 1 tsp of baking powder.

Further simplifying we have

9/4 tsp baking soda = 9 tsp baking powder

Hence we have 2.25 tsp baking soda = 9 tsp baking powder

2.25 can be written as 2 tsp + 1/4 tsp

Therefore we require 2.25tsp or 2+1/4 tsp of baking soda